C. What is the lenders APY?

thumbs up for correct solution

10. Anita borrows 540,000 at annual effective interest rate 3%. She repays this loan by paying off only the interest due at the end of each year to the lender and depositing a level amount Q at the end of each year into a sinking fund account paying 6% APY. The goal is to accumulate the full balance of the loan amount in the sinking fund at the end of 10 years. a. Find the level sinking fund deposit. b. What rate (AEIR) does Anita end up paying on this loan?
C. What is the lender's APY?
Answer #1

1.
=540000/(1.06^10-1)*6%=40968.69744

2.
Using financial calculator
N=10
PMT=540000*3%+40968.69744
PV=-540000
CPT I/Y=1.05%

3.
Lender APY=3%